OpenStudy (anonymous):

Find the limit as x approaches 0 (sin^2(3x))/(sqrt(121-9x^2)-11)

zepdrix (zepdrix):

Oh interesting problem! Ok ok lemme show you how this works.

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 0}\frac{\sin^2(3x)}{\sqrt{121-9x^2}-11}\]That looks correct so far?

OpenStudy (anonymous):

I can get 9x^2 as the denominator but I get stuck after that

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 0}\frac{\sin^2(3x)}{\sqrt{121-9x^2}-11}\color{royalblue}{\left(\frac{\sqrt{121-9x^2}+11}{\sqrt{121-9x^2}+11}\right)}\]You got through this step already? Ok good :) Which simplifies things down to,\[\large \lim_{x \rightarrow 0} \frac{\sin^2(3x)\left(\sqrt{121-9x^2}+11\right)}{-9x^2}\]Correct me if I'm wrong. I got a negative sign on the bottom. I think that's right.

OpenStudy (anonymous):

Sorry, I wrote the original equation wrong, it should be sqrt121+9x^2, so that would make it 9x^2

zepdrix (zepdrix):

ah ok that makes sense.

zepdrix (zepdrix):

Let's write this as the product of Limits. This might look a little strange, but go back to your limit laws to justify it. \[\large = \lim_{x \rightarrow 0}\frac{\sin^2(3x)}{9x^2}\cdot \left(\lim_{x \rightarrow 0}\sqrt{121-9x^2}+11\right)\]I think we're allowed to do this :) So from here, hmmm does the first limit look like a familiar identity? if not we can do a lil more to it.

zepdrix (zepdrix):

I'll just paste the idenity that I'm thinking of, so we're on the same page.\[\large \lim_{\theta \rightarrow 0}\frac{\sin \theta}{\theta}=1\]

zepdrix (zepdrix):

Woops i left a negative on the 9x^2, my bad. :p

OpenStudy (anonymous):

I follow you so far

OpenStudy (anonymous):

So we get sin3x/3x?

zepdrix (zepdrix):

Let's jimmy around with that first limit.\[\large \lim_{x \rightarrow 0}\frac{\sin^2(3x)}{9x^2} \quad =\quad \lim_{x \rightarrow 0}\frac{\sin^2(3x)}{(3x)^2} \quad = \quad \left(\lim_{x \rightarrow 0}\frac{\sin(3x)}{3x}\right)^2\] Yes very good. And following that rule, we can see that our \(\large \theta\) is 3x. So the whole first limit should simplify down to 1.

OpenStudy (anonymous):

Beautiful. That step was killing me

zepdrix (zepdrix):

Yay team. Imma throw it into wolfram real quick just to make sure I didn't make a boo boo anywhere.

zepdrix (zepdrix):

Cool! Yah looks like we did it correctly. And you understand what happens with the other limit part?

OpenStudy (anonymous):

Not sure.

OpenStudy (anonymous):

I assume we put 0 in and get 11 out?

OpenStudy (anonymous):

22 sorry

zepdrix (zepdrix):

Yes we can stick x=0 directly in since it's no longer a problem. 22? Yes very good.

OpenStudy (anonymous):

Awesome. I have been working on that all day. That one step was killing me. Thanks so much for your help

zepdrix (zepdrix):

no prob \c:/